OpenStudy (seattle12345):

can I have some help with this one? Factor completely 24x^2+30z-9.

Parth (parthkohli):

Looks like you can factor out the \(3\).

OpenStudy (seattle12345):

Ok so would it be 3(8z^2+10z-3)

Parth (parthkohli):

Yep, but it should be \(8x^2 \) instead of that z coming in.

OpenStudy (seattle12345):

oh yea sorry about that lol. so thats it?

OpenStudy (anonymous):

are you sure both x and z are in equation........?

OpenStudy (seattle12345):

so you get -24. two factors of -24 which add up to 10: 6+4... does it matter whether or not the factors are positive or negative? @nitz sorry that was a typo.

OpenStudy (anonymous):

wait ,,,,,,firstly what is the exact equation to be factorised???

OpenStudy (seattle12345):

24x^2+30x-9..... sorry about the confusion.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[3(8x ^{2}+10x-3)=3(8x ^{2}+12x-2x-3)=3{4x(2x+3)-1(2x+3)}\]

OpenStudy (anonymous):

sorry there is a bracket 3{4x(2x+3)-1(2x+3)} 3(4x-1)(2x+3)

OpenStudy (anonymous):

didnt get??????

OpenStudy (seattle12345):

no I got it, thanks for your help!